OpenStudy (anonymous):

Venus usually jogs to school and walks home. Today her rate going was 2 miles per hour more than her rate returning. If it took her 20 minutes to go to school and 30 minutes to return home by the same route, find her rate each way

OpenStudy (anonymous):

Didn't we already do this one?

OpenStudy (anonymous):

oh yes i just wanted to make sure the answer was correct. but how would you do 20/30(v)=d=30/60(v+2)? it comes out as a negative...

OpenStudy (anonymous):

\[ \frac{20}{60}(v)=d=\frac{30}{60}(v+2) \]

OpenStudy (anonymous):

You wrote it down wrong.

OpenStudy (anonymous):

Ooops my bad it should be: \[ \frac{30}{60}(v)=d=\frac{20}{60}(v+2) \]

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

This time it will be positive.

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

Okay if you want to think about it like a system of equations: \[\begin{array}{rcl} v_{go} &=& v_{come}+2 \\ 20\cdot v_{go} &=& 30\cdot v_{come} \end{array} \]

OpenStudy (anonymous):

Two variables, two equations.

OpenStudy (anonymous):

can't you just solve for v?